OpenStudy (anonymous):

Trig integration HELP PLEASE! step by step:( Ive been trying to solve this one for 2 hours now

OpenStudy (anonymous):

\[\int\limits_{?}^{?}\cos^4xsin^2x\]

OpenStudy (anonymous):

indefinite integral

OpenStudy (amistre64):

sin and cos are derivatives of each other, so there might be a way to think of this as:\[\int f'(g)g'~dx\to f(g)\]

OpenStudy (amistre64):

or using some creative trig identities

OpenStudy (amistre64):

but the even exponents might be screwing up the works alittle

hartnn (hartnn):

try writing sin^2 x as 1-cos^2 x then you'll have difference of two solvable integrals.

OpenStudy (anonymous):

I think youre supposed to use double angle identities but im not sure how.

OpenStudy (amistre64):

\[sin(a+a)=sin(a)cos(a)+sin(a)cos(a)\] \[sin(2a)=2sin(a)cos(a)\]

OpenStudy (anonymous):

This is how far I got \[\int\limits \frac{ 1 }{ 2 }(1-\cos2x)(\frac{ 1 }{ 2 } \sin2x)^2= \frac{ 1 }{ 8 } \int\limits \sin^2 (2x) -\sin^2 (2x)\cos(2x)\]